Grade 12th passPhysical Chemistry

Calculate mass of2.5moles of calcium if atomic mass of calcium is 40...

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Pawan Prajapati
5 Years ago
Explanation:

no of moles= given mass /atomic mass of the element

2.5=Mass/40

Mass=2.5×40=100g

Thanks